what is the interquatile range of the following data set 60 40 35 45 39
Answer by stanbon(75887)

what is the interquatile range of the following data set 60 40 35 45 39
---
Put the numbers in ascending order.
35, 39, 40, 45, 60
--
Find Q3, the data value that has 75% of the data below it:
locator: 0.75(5) = 15/4 = 3.75
Round up to 4:
Q3 is the 4th element = 45
----
Find Q1, the data value that has 25% of the data below it:
locator: 0.25(5) = 5/4 = 1.25
Round up to 2:
Q1 is the 2nd element = 39
-----
The interquartile range = Q3-Q1 = 45-39 = 6
==============================================
Cheers,
Stan H.
===============




http://www.google.com/search?hl=en&rls=com.microsoft:en-us:IE-SearchBox&rlz=1I7ADRA_en&defl=en&q=define:Interquartile+range&ei=NaXtS5WZDsHflge1laC1CA&sa=X&oi=glossary_definition&ct=title&ved=0CBIQkAE